### v1

@@ IV-B2 Pressure Data Measurement
Nevertheless, to compare both systems a qualitative comparison of plantar pressure distributions and force curves was conducted using visualizations from both systems for a representative trial (see Figure 8). The Zebris data represent averaged force curves for the three foot zones summarized from multiple gait cycles, whereas the orthosis data focus on one representative gait cycle from the same trial, which is also displayed in Figure 7. Although absolute force values are not directly comparable due to differences in sensor configuration, both systems show similar temporal trends across the three anatomical foot zones and in the overall force–time diagram. Compared to the Zebris treadmill, which provides only aggregated contact data, the SFO offers step-wise heatmap visualizations, enabling a detailed analysis of the internal plantar pressure distribution for each gait event (see Figure 7).
IV-C Smart Orthosis and Walking Aid: Perceived Haptic Feedback and User Experience
Shapiro–Wilk tests were conducted for each condition on the questionnaire items Noticeability, Non-Intrusiveness, Regular Usage, and Usefulness, and revealed that the data were not normally distributed. Therefore, an Aligned Rank Transform (ART) RM ANOVA was conducted on the four measures, revealing two significant main effects. A statistically significant main effect of Noticeability was found for Device with $F(1,21)=4.49$, $p=.046$, $\eta^{2}_{p}=.18$ (large effect size). However, post hoc pairwise comparisons using Wilcoxon signed-rank tests did not reveal a significant difference between Orthosis and Crutch ($p=.065$). Furthermore, a statistically significant main effect was found for Usefulness on Feedback Type with $F(1,21)=6.01$, $p=.023$, $\eta^{2}_{p}=.22$ (large effect size). Post hoc Wilcoxon signed-rank tests revealed a significant difference between Continuous and Pulsed Pattern ($p=.008$). Results of the RM ANOVA for all conditions are shown in Table III. Additionally, the overall results are presented as a boxplot with descriptive statistics in Figure 9.
TABLE III: ANOVA results for the four questionnaire items.
Measure Factor F Df p Sig.
Noticeability Device 4.49 1, 21 .046 *
Feedback 1.80 1, 21 .194
Device $\times$ Feedback 1.60 1, 21 .220
Non-Intrusiveness Device 0.06 1, 21 .815
Feedback 0.50 1, 21 .487
Device $\times$ Feedback 2.24 1, 21 .150
Regular Usage Device 0.00 1, 21 .966
Feedback 2.57 1, 21 .124
Device $\times$ Feedback 0.77 1, 21 .390
Usefulness Device 0.00 1, 21 1
Feedback 6.01 1, 21 .023 *
Device $\times$ Feedback 0.00 1, 21 1
Figure 9: Boxplots show agreement ratings for four measures on haptic feedback via the foot orthosis and the forearm crutch, ranging from 1 (strongly disagree) to 7 (strongly agree). It displays the interquartile range (box), 1.5×IQR range (whiskers), median (line), and mean (black dot). Mean and standard deviation (±) values are shown below each boxplot.
IV-C1 Correlation of Ratings Across Devices
To further investigate consistency in user perception across devices, a Spearman rank correlation was conducted between ratings for the same Feedback Type on the Orthosis and the Crutch. Results showed a significant positive correlation for Noticeability in the Pulsed Pattern feedback condition ($r_{s}=.86$, $p=.006$), suggesting a high level of consistency in perceived noticeability across both devices. Furthermore, in the Continuous feedback condition, a significant positive correlation was found for Regular Usage ($r_{s}=.78$, $p=.023$), while the remaining conditions showed positive trends, but did not reach statistical significance: Noticeability ($r_{s}=0.62$, $p=.099$), Non-Intrusiveness ($r_{s}=0.56$, $p=.152$), and Usefulness ($r_{s}=0.58$, $p=.133$).
IV-C2 Post-Experiment Questionnaire
IV-D Smartphone Application: Usability Evaluation
The mean SUS score was 80.62 (SD = 10.59), as shown in the boxplot diagram (Figure 10(a)). According to the interpretation guidelines by Bangor et al. [110], this score falls within the “Acceptable” range based on the Acceptability Ranges and is considered “Good” according to the Adjective Ratings, indicating a high level of usability. Ratings for the ten individual SUS items are visualized in the radar plot (Figure 10(b)), with values ranging from 6.56 to 9.38 (normalized on a 0–10 scale). Furthermore, the raw ratings of the subscales are summarized in Table IV using descriptive statistics.
TABLE IV: Descriptive statistics of the SUS items (M = Mean, SD = Standard Deviation).
@@ IV-D Smartphone Application: Usability Evaluation
(a) Boxplot - SUS score
(b) Radar plot - 10 items
Figure 10: System Usability Scale (SUS) results: a) Boxplot showing the overall SUS score; b) Radar plot visualizing the ten individual item scores, normalized on a 0–10 scale. Asterisks (*) indicate reverse-scored items.
V Discussion
We developed a smart insole concept for monitoring plantar pressure and gait activity within an orthopedic foot orthosis. The system provides haptic feedback through either the orthosis or a corresponding forearm crutch, and enables remote monitoring via a smartphone app. In an experimental user study with eight participants, we demonstrated reliable detection of gait phases using force sensors, as well as accurate temporal identification of gait events. Furthermore, our results indicate a preference for vibration pulse–based haptic feedback, delivered either inside the orthosis or via the crutch. In addition, the corresponding smartphone app was rated as highly user-friendly, which underscores the general applicability of the system. However, the validation also revealed limitations of the data processing as spatial parameters were not measured with sufficient precision, indicating the need for further refinement of the sensing approach. Moreover, the lack of statistically significant differences in the haptic feedback testing may be due to the small sample size (n = 8), which limits the statistical power. Overall, these findings highlight both the feasibility and potential, while also revealing opportunities for further research and improvement, which are discussed in the following subsections.
V-A Analysis of Gait Data Accuracy (RQ1)
RQ1 focused on evaluating the accuracy of plantar pressure and gait measurements obtained from the developed SFO compared to a clinical reference system.
Our findings showed that the recorded temporal parameters for the swing and cycle phases indicated significantly shorter durations when measured with the SFO compared to the treadmill. However, cycle duration showed a strong correlation between both systems, whereas swing duration did not reach statistical significance ($p=.067$). Overall, swing duration showed the lowest accuracy among all temporal parameters, with only 60% accuracy and notably higher standard deviations. This may be explained by the observation that some FSR sensors were constantly triggered and did not return to zero during offloading, which could have impaired the correct detection of events from stance to swing. Although the stance phase showed no significant differences and achieved high accuracy of 96%, the correlation between both systems was not significant. These inconsistencies are likely caused by differences in event detection, as the SFO relies on FSR-based thresholds, whereas the treadmill reference system uses ground reaction forces. Additionally, some triggering issues of certain FSR sensors may have introduced inconsistent deviations in individual measurements, which could explain the reduced correlation, even though the average stance durations between both systems were similar. As highlighted in previous research, lower walking speeds (0.8–1.0 m/s) compared to moderate walking speeds (1.2–1.4 m/s) are associated with higher error rates in temporal parameters (e.g., stance and swing time), as well as spatial parameters (e.g., stride length) [111]. As the walking speed in our experiment was relatively low (0.417 m/s) to ensure participant comfort while walking with the orthosis, this factor may have contributed to reduced accuracy in gait parameter estimation.
The measured accuracy of spatial cycle and step length showed that it is possible to achieve an accuracy above 80%. Overall, the estimation of spatial parameters based on a single smart insole systematically underestimated step length and cycle length compared to treadmill-based measurements, however, it showed positive correlation trends across trials. One explanation for these systematically lower values is that stride length is only estimated, as the insole was worn unilaterally, while the other foot was not equipped with an insole to provide additional data. By identifying the systematic offset, a post-processing correction factor could be applied to improve the dataset for further analyses.
The evaluation of pressure data showed that gait events can be reliably detected and continuous pressure distribution can be measured during both walking and standing. Although the comparison with the treadmill was performed on a qualitative basis, the results showed that the three foot zones (forefoot, midfoot, and hindfoot) exhibited similar temporal patterns. This highlights the added value of the SFO, as it enables step-wise heatmap visualizations and detailed insights into the internal plantar pressure distribution without post-processing, which are not achievable with stationary treadmill systems.
In summary, our results confirm previous findings showing that a single IMU is sufficient for gait monitoring [47] and that the recommended configuration of 15 FSRs [42] is suitable for reliable pressure measurement. Overall, our findings indicate that using a single insole combined with targeted data pre-processing can be sufficient for gait and pressure monitoring. However, further refinements are recommended to improve measurement accuracy and enhance data reliability.
@@ V-A Analysis of Gait Data Accuracy (RQ1)
V-B User Perception of Haptic Feedback (RQ2)
RQ2 investigated how different haptic feedback types (Continuous vs. Pulsed Pattern) implemented in an orthosis and a walking aid affect user perception regarding Noticeability, Non-Intrusiveness, Regular Usage, and Usefulness. Descriptive results showed that participants generally rated the Pulsed Pattern feedback higher than the Continuous feedback across all four items for both devices. An exception was found for Non-Intrusiveness, where the Pulsed Pattern received lower ratings specifically in the orthosis condition.
Inferential analysis using an ART RM ANOVA revealed a significant main effect of Device on Noticeability, as the crutch was rated significantly higher than the orthosis. One possible explanation is that some participants had reduced sensation in the lower limbs, which made the haptic signal (especially Continuous feedback) less perceptible when delivered via the foot. Although previous work has suggested that foot-based feedback may be preferable to hand-based in some applications [112], in this specific use case the hand appears to be more sensitive to the vibration signal using one ERM actuator. For Non-Intrusiveness, the ratings varied more strongly across conditions. In particular, the Continuous feedback was rated lower than the Pulsed Pattern when applied via the crutch, which indicates that continuous vibrations may be perceived as more intrusive in this form factor. The Regular Usage showed similar ratings, however, without significances. Ratings for perceived Usefulness showed a significant effect of feedback type, indicating that the Pulsed Pattern was perceived as more useful than the Continuous feedback. This result aligns with the post-study feedback, in which most participants expressed a clear preference for the pulsed signal. Although other trends were observed, they did not reach statistical significance, which is probably due to the small size and limited statistical power of the study.
These findings support previous research on the utility of ERM-based vibration feedback [113]. While a single ERM actuator ensured sufficient noticeability in the crutch, we recommend using at least two actuators inside the orthosis to enhance perceptibility. Overall, our results underscore the importance of tailoring haptic feedback implementation to the device form factor and user preferences. In line with prior work emphasizing the need to evaluate haptic feedback in assistive systems [75], our study contributes initial insights into user perception, acceptance, and effectiveness.
V-C Evaluation of Smartphone App Usability (RQ3)
@@
Regarding RQ3, which addressed the usability of the accompanying mobile application for visualizing gait and pressure data, our results showed an acceptable level of usability. The SUS evaluation reached an average score of 80.62 (SD = 10.59), which is comparable to the results reported by Resch et al. [65] for a smart insole app (type: basic), with a mean SUS score of 77.5 (SD = 16.74). In comparison to other digital health applications, our findings also exceed the average usability benchmark identified by Hyzy et al. [114], who conducted a meta-analysis and reported a mean SUS score of 68.05 (SD = 14.05) across various digital health apps.
@@
Overall, the subscale ratings revealed that item 4 (need for support) and item 7 (learnability) received the highest scores. This indicates that participants largely agreed with the statement ”most people would learn to use this app very quickly” (item 7) and disagreed with the statement ”I would need the support of a technical person to use the system” (item 4), which was reverse-scored. In contrast, item 1 (frequency of use) received the lowest average rating (M = 3.62, SD = 0.92 on a 5-point Likert scale), which shows that participants potentially were less willing to use the app frequently. This might be due to the fact that some participants were only temporarily affected by foot-related conditions and therefore did not perceive frequent use of the app as necessary. However, to address the potential limited interest in frequent use, future iterations of the app could incorporate targeted strategies to increase user engagement. For example, personalized notifications and reminders could encourage consistent daily use, particularly for self-monitoring tasks such as the integrated pain-level diary. Previous research suggested that gamification elements in health interventions (e.g., progress tracking or reward systems) can have positive effects on user experience [115]. Therefore, the integration of personalized measurable goals and feedback could support user motivation and foster continued use, especially in long-term rehabilitation contexts.
V-D Implications
This work contributes to the broader field of assistive gait monitoring and feedback interaction technologies in rehabilitation and preventive healthcare. First, the smart orthosis demonstrated the feasibility of accurate pressure and motion sensing for mobile gait analysis. This enables continuous monitoring of plantar pressure distributions using heatmaps, which is not possible with conventional systems. Such detailed information may help identify potential overloads early and complement laboratory-based gait analysis. Second, the evaluation of haptic feedback confirmed that pattern-based feedback is generally perceived as more useful for alert notifications, particularly in cases of pressure overload. However, noticeability varied substantially between the orthosis and crutch, highlighting the importance of customizable feedback modules that adapt to user preferences and functional needs. Third, the development of the mobile application underlines the importance of user-centered feedback design. Future applications could integrate features such as gamified elements or AI-based chatbots to support daily engagement and simplify the interpretation of health data. These findings can guide future research and the development of intelligent footwear, multimodal feedback solutions, and interactive mobile health applications, contributing to the fields of human-computer interaction (HCI) and biomedical engineering.
V-E Limitations

### v2

@@ 4.1.2 Pressure Data Measurement
4.2 Application Usability: Evaluation of Smartphone App Usability
The mean SUS score was 80.62 (SD = 10.59), as shown in the boxplot diagram (Figure 9a). According to the interpretation guidelines by Bangor et al. Bangor et al. (2009), this score falls within the “Acceptable” range based on the Acceptability Ranges and is considered “Good” according to the Adjective Ratings, indicating a high level of usability. Ratings for the ten individual SUS items are visualized in the radar plot (Figure 9b), with values ranging from 6.56 to 9.38 (normalized on a 0–10 scale). Furthermore, the raw ratings of the subscales are summarized in Table 3 using descriptive statistics.
((a))
((b))
Figure 9: (a) Boxplot of overall SUS scores and the corresponding acceptability categories; (b) Radar plot visualizing the ten individual item scores, normalized on a 0–10 scale. Asterisks (*) indicate reverse-scored items.
Table 3: Descriptive statistics of the SUS items (M = Mean, SD = Standard Deviation).
@@ 4.2 Application Usability: Evaluation of Smartphone App Usability
4.3 Feedback Evaluation: User Perception of Haptic Feedback (Smart Orthosis and Walking Aid)
Shapiro–Wilk tests were conducted for each condition on the questionnaire items Noticeability, Non-Intrusiveness, Regular Usage, and Usefulness, and revealed that the data were not normally distributed. Therefore, an Aligned Rank Transform (ART) RM ANOVA was conducted on the four measures, revealing two significant main effects. A statistically significant main effect of Noticeability was found for Device with $F(1,21)=4.49$, $p=0.046$, $\eta^{2}_{p}=0.18$ (large effect size). Post hoc pairwise comparisons using Wilcoxon signed-rank tests did not reveal a significant difference between Orthosis and Crutch ($p=0.065$). However, the corresponding effect size ($r=0.483$, moderate-to-large) suggests a potentially meaningful difference that may not have reached significance due to the limited sample size and resulting low statistical power. Furthermore, a statistically significant main effect was found for Usefulness on Feedback Type with $F(1,21)=6.01$, $p=0.023$, $\eta^{2}_{p}=0.22$ (large effect size). Post hoc Wilcoxon signed-rank tests revealed a significant difference between Continuous and Pulsed Pattern ($p=0.008$).
Results of the ART RM ANOVA and post hoc pairwise comparisons for all conditions are shown in Table 4. Additionally, the overall results are presented as a boxplot with descriptive statistics in Figure 10.
Figure 10: Boxplots show agreement ratings for four measures on haptic feedback via the foot orthosis and the forearm crutch, ranging from 1 (strongly disagree) to 7 (strongly agree). It displays the interquartile range (box), 1.5 × IQR range (whiskers), median (line), and mean (black dot). Mean and standard deviation (±) values are shown below each boxplot. Asterisks (*) indicate statistically significant results ($p<0.05$).
Table 4: Results of the ART RM ANOVA and post-hoc Wilcoxon signed-rank tests for the four questionnaire items.
ART RM ANOVA Pairwise Comparisons (Wilcoxon)
[…]
Noticeability Device 4.49 1, 21 0.046 * 0.176 25.0 0.065 0.483 moderate
Feedback 1.80 1, 21 0.194 0.079 22.0 0.187 0.310 moderate
Device × Feedback 1.60 1, 21 0.220 0.071
Non-Intrusiveness Device 0.06 1, 21 0.815 0.003 27.5 0.653 0.106 small
Feedback 0.50 1, 21 0.487 0.023 42.5 0.811 0.046 small
Device × Feedback 2.24 1, 21 0.150 0.096
Regular Usage Device 0.00 1, 21 0.966 0.000 42.5 0.807 0.099 small
Feedback 2.57 1, 21 0.124 0.109 50.0 0.129 0.326 moderate
Device × Feedback 0.77 1, 21 0.390 0.035
Usefulness Device 0.00 1, 21 1.000 0.000 24.5 0.851 0.076 small
Feedback 6.01 1, 21 0.023 * 0.220 36.0 0.008 0.703 large
Device × Feedback 0.00 1, 21 1.000 0.000
Note. Asterisks (*) indicate statistically significant results ($p<0.05$). Effect sizes are reported as partial $\eta^{2}$ for ART RM ANOVA and as rank-biserial correlation ($r$) for Wilcoxon signed-rank tests. Magnitude classification of $r$: small ($r<0.3$), moderate ($0.3\leq r<0.5$), large ($r\geq 0.5$Cohen (2013).
4.3.1 Correlation of Ratings Across Devices
To further investigate consistency in user perception across devices, a Spearman rank correlation was conducted between ratings for the same Feedback Type on the Orthosis and the Crutch. Results showed a significant positive correlation for Noticeability in the Pulsed Pattern feedback condition ($r_{s}=0.86$, $p=0.006$), suggesting a high level of consistency in perceived noticeability across both devices. Furthermore, in the Continuous feedback condition, a significant positive correlation was found for Regular Usage ($r_{s}=0.78$, $p=0.023$), while the remaining conditions showed positive trends, but did not reach statistical significance: Noticeability ($r_{s}=0.62$, $p=0.099$), Non-Intrusiveness ($r_{s}=0.56$, $p=0.152$), and Usefulness ($r_{s}=0.58$, $p=0.133$).
4.3.2 Post-Experiment Questionnaire
5 Discussion
We developed a sensorized system for monitoring plantar pressure and gait activity within an orthopedic foot orthosis. The system provides haptic feedback through either the orthosis or a corresponding forearm crutch and enables remote monitoring via a smartphone application. In an experimental user study with eight participants, we evaluated the system’s feasibility and investigated the influence of device location and feedback modality on user perception of vibrotactile biofeedback. The findings confirm the system’s technical feasibility for gait phase detection, the usability of the accompanying mobile application, and provide new insights into haptic feedback perception. The following subsections discuss the results in detail, including implications, limitations, and directions for future research.
5.1 System Feasibility and Application Usability
5.1.1 System Feasibility: Accuracy of Plantar Pressure and Gait Parameters (SRQ1)
SRQ1 focused on evaluating the accuracy of plantar pressure and gait measurements obtained from the developed SFO compared to a clinical reference system. Our findings showed that the recorded temporal parameters for the swing and cycle phases indicated significantly shorter durations when measured with the SFO compared to the treadmill. However, cycle duration showed a strong correlation between both systems, whereas swing duration did not reach statistical significance ($p=0.067$). Overall, swing duration showed the lowest accuracy among all temporal parameters, with only 60% accuracy and notably higher standard deviations. This may be explained by the observation that some FSR sensors were constantly triggered and did not return to zero during offloading, which could have impaired the correct detection of events from stance to swing. Although the stance phase showed no significant differences and achieved high accuracy of 96%, the correlation between both systems was not significant. These inconsistencies are likely caused by differences in event detection, as the SFO relies on FSR-based thresholds, whereas the treadmill reference system uses ground reaction forces. Additionally, some triggering issues of certain FSR sensors may have introduced inconsistent deviations in individual measurements, which could explain the reduced correlation, even though the average stance durations between both systems were similar. As highlighted in previous research, lower walking speeds (0.8–1.0 m/s) compared to moderate walking speeds (1.2–1.4 m/s) are associated with higher error rates in temporal parameters (e.g., stance and swing time), as well as spatial parameters (e.g., stride length) Ganguly et al. (2023). As the walking speed in our experiment was relatively low (0.417 m/s) to ensure participant comfort while walking with the orthosis, this factor may have contributed to reduced accuracy in gait parameter estimation.
The measured accuracy of the spatial parameter (cycle length) indicated that an accuracy above 83% can be achieved. However, the estimation of spatial gait parameters using a single smart insole systematically underestimated cycle length compared to treadmill-based reference measurements, and no significant correlation was found between both measurement systems. One explanation for these systematically lower values is that the insole was worn unilaterally, while the other foot was not equipped with a sensorized insole to provide complementary data. Furthermore, wrong triggering of the algorithm could have led to incorrect gait event detection. By identifying this systematic offset, a post-processing correction factor could be applied to improve the dataset for further analyses.
The evaluation of pressure data showed that gait events can be reliably detected and continuous pressure distribution can be measured during both walking and standing. Although the comparison with the treadmill was performed on a qualitative basis, the results showed that the three foot zones (forefoot, midfoot, and hindfoot) exhibited similar temporal patterns. This highlights the added value of the SFO, as it enables step-wise heatmap visualizations and detailed insights into the internal plantar pressure distribution without post-processing, which are not achievable with stationary treadmill systems. In summary, our results confirm previous findings showing that one IMU is sufficient for gait monitoring Sabatini et al. (2005) and that the recommended configuration of 15 FSRs Razak et al. (2012) is suitable for reliable pressure measurement. However, further refinements are recommended to improve measurement accuracy and enhance data reliability.
Compared to related studies on smart insoles and gait monitoring systems, the accuracy levels achieved in this work are within a comparable range. For example, D’arco et al. D’arco et al. (2024) reported high performance in activity recognition using smart insoles, with accuracies ranging from 70% to 99%. This demonstrates that comparable performance levels can be achieved even without additional post-processing. Despite differences in sensor configuration (type, quantity, and placement), as well as varying application contexts, testing scenarios and data processing, the present study demonstrates that reliable gait analysis can also be achieved using a single insole. With further iterative refinements of the sensing and calibration procedures, even higher accuracy levels may be achieved in future implementations.
5.1.2 Application Usability: Evaluation of Smartphone App Usability (SRQ2)
@@
Regarding SRQ2, which addressed the usability of the accompanying mobile application for visualizing gait and pressure data, our results showed an acceptable level of usability. The SUS evaluation reached an average score of 80.62 (SD = 10.59), which is comparable to the results reported by Resch et al. Resch et al. (2024) for a smart insole app (type: basic), with a mean SUS score of 77.5 (SD = 16.74). In comparison to other digital health applications, our findings also exceed the average usability benchmark identified by Hyzy et al. Hyzy et al. (2022), who conducted a meta-analysis and reported a mean SUS score of 68.05 (SD = 14.05) across various digital health apps.
@@
Overall, the subscale ratings revealed that item 4 (need for support) and item 7 (learnability) received the highest scores. This indicates that participants largely agreed with the statement “most people would learn to use this app very quickly” (item 7) and disagreed with the statement “I would need the support of a technical person to use the system” (item 4), which was reverse-scored. In contrast, item 1 (frequency of use) received the lowest average rating (M = 3.62, SD = 0.92 on a 5-point Likert scale), which shows that participants potentially were less willing to use the app frequently. This might be due to the fact that some participants were only temporarily affected by foot-related conditions and therefore did not perceive frequent use of the app as necessary. However, to address the potential limited interest in frequent use, future iterations of the app could incorporate targeted strategies to increase user engagement. For example, personalized notifications and reminders could encourage consistent daily use, particularly for self-monitoring tasks such as the integrated pain-level diary. Previous research suggested that gamification elements in health interventions (e.g., progress tracking or reward systems) can have positive effects on user experience Alzghoul (2024). Therefore, the integration of personalized measurable goals and feedback could support user motivation and foster continued use, especially in long-term rehabilitation contexts.
5.2 Feedback Evaluation: User Perception of Haptic Feedback (MRQ)
The MRQ investigated how different haptic feedback types (Continuous vs. Pulsed Pattern) implemented in an orthosis and a walking aid affect user perception regarding Noticeability, Non-Intrusiveness, Regular Usage, and Usefulness. Descriptive results showed that participants generally rated the Pulsed Pattern feedback higher than the Continuous feedback across all four items for both devices. A likely reason is that the rhythmic structure of the pulsed signal was perceived as more distinct and meaningful, making it easier to recognize and less monotonous compared to a continuous vibration. An exception was found for Non-Intrusiveness, where the Pulsed Pattern received lower ratings specifically in the orthosis condition.
Inferential analysis using an ART RM ANOVA revealed a significant main effect of Device on Noticeability, as the crutch was rated significantly higher than the orthosis. One possible explanation is that some participants had reduced sensation in the lower limbs, which made the haptic signal (especially Continuous feedback) less perceptible when delivered via the foot. Although previous work has suggested that foot-based feedback may be preferable to hand-based in some applications Gurari et al. (2009), in this specific use case the hand appears to be more sensitive to the vibration signal using one ERM actuator. This highlights that feedback effectiveness is not only dependent on signal type, but also on feedback location and tactile sensitivity at different body sites and devices.
For Non-Intrusiveness, the ratings varied more strongly across conditions. In particular, the Continuous feedback was rated lower than the Pulsed Pattern when applied via the crutch, which indicates that continuous vibrations may be perceived as more intrusive in this form factor. The Regular Usage showed similar ratings, however, without significance. Ratings for perceived Usefulness showed a significant effect of feedback type, indicating that the Pulsed Pattern was perceived as more useful than the Continuous feedback. These findings highlight that temporally modulated vibration signals are more easily interpreted and less annoying than continuous signals in wearable biofeedback applications. This result aligns with the post-study feedback, in which most participants expressed a clear preference for the pulsed signal. Although other trends were observed, they did not reach significance, which is probably due to the small size and limited statistical power of the study.
These findings support previous research on the utility of ERM-based vibration feedback Choi and Kuchenbecker (2013). However, our results extend this work by demonstrating that both pattern and device location influence the perception of vibrotactile feedback in orthopedic assistive devices. While a single ERM actuator ensured sufficient noticeability in the crutch, we recommend using at least two actuators inside the orthosis to enhance perceptibility. Overall, our results underscore the importance of tailoring haptic feedback implementation to the device form factor and user preferences. Therefore, designers should consider not only the feedback modality but also where and how it is delivered, and ensure that the signal is both interpretable and acceptable during everyday use. In line with prior work emphasizing the need to evaluate haptic feedback in assistive systems Resch et al. (2025), our study contributes initial insights into user perception, acceptance, and effectiveness.
5.3 Implications
The presented system was designed as a general-purpose gait monitoring and feedback platform that can be applied both to individuals with foot conditions and for preventive or rehabilitative use in broader populations. While no clinical validation has yet been conducted, the system’s sensing and feedback capabilities hold potential for future application, particularly for monitoring plantar pressure distribution and detecting overload patterns during daily activities.
The primary contribution of this work lies in advancing the understanding of how vibration-based feedback can be applied and perceived in foot-related assistive systems. While the developed hardware and software components served as necessary groundwork, they were mainly used for system verification. Building on this foundation, the central contribution is the investigation of real-time haptic overload notifications driven by the integration of plantar pressure and motion measurements. The fused dataset from pressure and motion sensors enabled a closed-loop feedback mechanism that responds to gait events. Our results demonstrate that such multimodal data can be feasibly integrated to trigger vibration cues. Based on this, the evaluation of haptic feedback provided new insights into the perception and preferences related to vibration-based notifications in assistive systems. The findings show that pattern-based feedback is perceived as more effective for overload alerts, while perceptual differences between the orthosis and the walking aid highlight the need for differentiated actuator placement, intensity modulation, and an increased number of actuators to ensure sufficient noticeability. These insights contribute to a clearer understanding of multimodal feedback perception in foot augmentation systems and can inform design guidelines for future adaptive vibration strategies in assistive devices.
Overall, these findings extend existing knowledge on isolated prototype developments by combining system integration, technical validation, and user-centered evaluation. These implications can guide future research and the development of intelligent footwear, multimodal feedback solutions, and interactive mobile health applications, contributing to the fields of HCI and biomedical engineering.
5.4 Limitations